An employment agency specializing in temporary construction help pays heavy equipment operators $ 132 per day and general laborers $81 per day. If thirty-two people were hired and the payroll was $3510, how many heavy equipment operators were​ employed? How many​ laborers?
Answer by mananth(16946)


Let number of heavy equipment operators be x
and laborers be y
Thirty-two people were hired

x+y =32
Heavy equipment operators $ 132 per day
and general laborers $81 per day
the payroll was $3510,
132x + 81y = 3510
x Heavy equipment operators $ 132
y labourers $ 81


1.00 x + 1.00 y = 32.00


132.00 x + 81.00 y = 3510.00 .............2
Eliminate y
multiply (1)by -81.00
Multiply (2) by 1.00
-81.00 x -81.00 y = -2592.00
132.00 x 81.00 y = 3510.00
Add the two equations
51.00 x = 918.00
/ 51.00
x = 18.00
plug value of x in (1)
1.00 x + 1.00 y = 32.00
18.00 + 1.00 y = 32.00
1.00 y = 14.00
y = 14.00
Ans x = 18.00
y = 14.00
18 Heavy equipment operators
14 labourers



how many heavy equipment operators were​ employed? How many​ laborers?